# Missing Money and Market-Based Adequacy in Deeply Decarbonized Power Systems with Long-Duration Energy Storage

### v1

@@ II-A4 Market clearing conditions
Finally, agent-specific decisions must comply with system-wide constraints. In particular, market-clearing conditions ensure a balance between supply and demand for both energy and CMs, and a constraint on the annual expected emissions ensures that the CO2 emissions from electricity generation do not exceed the cap. The dual variables of these constraints, $\lambda^{E}_{\omega,t}$, $\lambda^{C}$, and $\lambda^{CT}$, can be interpreted as market prices (weighted with scenario probabilities and time interval durations).
@@
$=d_{\omega t}^{\text{fix}}-d_{\omega t}^{\text{flex}}\quad\forall t,\omega;\hskip 0.0pt(\lambda^{E}_{t,\omega})$ (8a)
$\sum_{g}c^{CM}_{g}+\sum_{s}c^{CM}_{s}=d^{CM};\hskip 0.0pt(\lambda^{CM})$ (8b)
$\sum_{g,t,\omega}w_{\omega t}\delta_{\omega}q_{\omega t,g}EF_{g}\leq\bar{E}^{CT};\hskip 0.0pt(\lambda^{CT})$ (8c)
@@ II-B Solution strategy
The market players’ problems are convex quadratic with linear constraints, making their Karush–Kuhn–Tucker (KKT) conditions necessary and sufficient for optimality [28]. By combining KKT conditions and applying the integrability principle of variational inequalities [29], the solution to the central planner’s convex quadratic optimization problem equates to the equilibrium solution [29].
@@ II-C Capacity accreditation
We estimate capacity credits using the marginal accreditation approach, supported by academic literature and increasingly adopted in industry [15]. Credits are calculated for each generator $g\in\mathcal{G}$ and storage unit $s\in\mathcal{S}$, with storage evaluated across various durations $\zeta\in\mathcal{Z}$. Thus, the resource set for accreditation is $\mathcal{R}=\mathcal{G}\cup(\mathcal{S}\times\mathcal{Z})$.
The accreditation methodology follows similar principles to those proposed by the NESO for marginal EFC [19] and builds upon the structure presented by Wang et al. [21] to consider multiple storage durations. We first determine the baseline EUE resulting from dispatching the welfare-maximising capacity mix under the price-capped market. Unlike in conventional models with perfectly inelastic demand, the definition of unserved energy is less straightforward when demand is price-elastic [^4] It is also worth noting that the solution to our equilibrium problem, while maximising social welfare, may not be EUE minimising. EUE-minimising storage dispatch is one of the assumptions in Zachary et al. [15] for integrating ES into capacity markets, which may not hold in our framework, as the storage dispatch is driven by economics.. In this work, we follow an approach similar to [30]. Considering all demand with WTP greater than or equal to the price cap as inelastic, we record any shortfall of consumption with respect to that demand as unserved [^5] Note that the PC not only limits the energy price levels during scarcity but also removes part of the flexible demand, which would voluntarily adjust its level based on the energy price.. The unserved energy $l_{\omega t}$ is given by:
@@
Subsequently, we perturb the capacity mix by adding a hypothetical perfect generator with small capacity $\epsilon$ and re-solve the model to obtain the reference EUE. Then, for each resource, including storage with specific durations, we add $\epsilon$ capacity of this resource to the mix and calculate the capacity credit as the ratio of the EUE reduction to that of the perfect generator. The complete process is presented in Algorithm 1.
@@
  Output: Capacity credit $CC_{r},\forall r\in R$
@@
  Fix $c_{r}\leftarrow\hat{c}_{r},\forall r\in R$
@@
   Compute capacity credit: $CC_{r}=\frac{\text{EUE}_{0}-\text{EUE}_{r}}{\text{EUE}_{0}-\text{EUE}_{\text{ref}}}$
@@ II-D Capacity market parameters
We construct the capacity demand curve by determining the capacity target and the corresponding WTP function. To determine the WTP, we rely on the system-wide net CONE and an assumed elasticity of demand for capacity. Following [31], we calculate the capacity target $CT$ using the target technology-specific capacity installations and the estimated capacity credits.
$CT=\sum_{g}CC_{g}\dot{c}_{g}+\sum_{s}f^{CC}(\dot{\zeta}_{s})\dot{c}_{s}^{P}$ (11)
Net CONE is defined as the difference between a technology’s annualised fixed and operational costs and its expected net revenues from the energy market, including ancillary services where applicable. Traditionally, CMs estimate net CONE for a reference technology, typically a flexible conventional generator whose accredited capacity is based on its availability. We follow industry practice and use CCGT as the reference technology, assuming, for demonstration purposes, its perfect availability and flexibility.
Assuming that the price cap distorts prices only during hours with scarce supply and that the CC captures the impact of the price cap distortion during those hours on the financial bottom line of a resource, the net CONE per accredited MW in a long-term equilibrium should theoretically be the same for all resources [32]. Anticipating this long-run outcome, we estimate alternative values for CC, hereafter referred to as required CC, with a clear economic interpretation [33]. Eq. (12) computes the required $CC^{\text{req}}_{r}$ as the ratio of the net CONE per installed MW of technology $r$ ($\text{NC}^{\text{uc}}_{r}$) divided by the net CONE per installed MW of the reference technology ($\text{NC}_{\text{ref}})$666 The required CC is valid only for resources that have positive installed capacity in the social-welfare maximizing capacity mix. Only these resources should recover their costs from the capacity and energy markets in the long run.. Eq. (12) assumes that the accredited capacity of the reference technology is equal to its installed capacity.
@@
$\text{NC}_{\text{ref}}=\frac{\text{NC}^{\text{uc}}_{r}}{CC^{\text{req}}_{r}}\rightarrow CC_{r}^{\text{req}}=\frac{\text{NC}^{\text{uc}}_{r}}{\text{NC}_{\text{ref}}}\quad\{\forall r\in\mathcal{R}|\hat{c}_{r}>0\}$ (12)
@@ II-E Experiment design
For our simulations, we use the model presented in II-A in two distinct market structures: an energy-only market (EOM), where CM-related variables and constraints are omitted, and a combined energy and CM (E+CM). Each structure can be run with different energy demand functions, allowing the price to spike to VOLL or be capped at the PC, set at an administratively defined level significantly lower than VOLL. When prices are allowed to spike to VOLL, the amount of the price-elastic demand is $D^{\text{flex}}_{\omega,t}$. With the PC, the width of the linear part of the demand curve reduces by $D^{\text{flex}}_{\omega,t}\cdot\frac{\text{VOLL}-\text{PC}}{\text{VOLL}}$ and the flat part of the demand curve increases in width by the same amount.
Effectively, we run the following main combinations: EOM VOLL is an energy-only market WTP function rising to VOLL; EOM PC is an energy-only market with WTP capped at PC; E+CM is an energy+capacity market with WTP capped at PC. For comparative and accreditation purposes, we also run a case denoted EOM PC opt. mix, where the price is capped, but the capacity mix is fixed to the welfare-maximising mix from EOM VOLL.
The welfare-maximising capacity mix is determined by solving the EOM VOLL model. Next, the EOM PC opt. mix model provides us with the baseline level for $\text{EUE}_{o}$, which is used to calculate capacity credits. Note that when the price cap is applied, the welfare-maximizing mix is no longer economically feasible due to limited price spikes during scarcity, which curb the cost recovery. Finally, to identify financially viable capacity mixes in the presence of a cap on energy prices, we solve the E+CM and EOM PC models, which include and lack a capacity market, respectively.
@@ II-F Performance metrics
We use two primary performance metrics to evaluate the effectiveness of the E+CM model. The first metric is the social welfare, which is calculated as the difference between the consumer benefit (6a) and capital and operating costs. By construction, the SW of the EOM VOLL run will be the highest because in that case, the objective function of the optimisation problem is identical to the SW metric. The SW lost in the other runs measures the cost of a price cap or the effectiveness of the E+CM design in terms of SW. The second metric is EUE, as defined earlier in (10). In addition to our main metrics, we discuss how different the E+CM capacity mix is compared to the welfare-maximising capacity mix.
@@
In this methodological framework, three factors can contribute to social welfare inefficiencies in the E+CM market. First, calibrating CM parameters is challenging, and estimation errors in capacity credits or net-CONEs can impact the results. Second, imposing price caps narrows price spreads, disrupting storage operational patterns. Third, the unserved demand may be inefficiently allocated between periods as the model assigns a constant value of PC to every MWh unserved. We evaluate SW according to the true WTP function, following Joskow and Tirole [34] in assuming efficient (non-price) rationing so that customers with the highest WTP are prioritised, but this means the average per-MWh cost of unserved energy rises with the amount unserved in that period. Under the alternative assumption of random rationing, the welfare cost of every MWh unserved would be equal, but the total cost of any given reduction in load would be higher than if it was targeted at the lower-WTP customers.
@@
The latter two effects above relate solely to the introduction of a PC, and can be assessed by running the EOM PC case with optimal mix obtained from EOM VOLL (EOM PC opt. mix. mentioned earlier). To isolate these effects, we need to obtain the optimal true WTP-aware discharge schedule for the storage and remove the effects of true WTP inobservability on the EUE distribution, i.e., isolating the third effect. First, we fix the generation and charging decisions resulting from the original run. Subsequently, we replace the price-capped WTP with the true WTP, allowing it to reach the VOLL. Finally, we redispatch the mix, with only the discharging variables unfixed, which will now be scheduled to maximise true SW. Because the charging decisions of the storage are fixed, the schedule will only change discharging to better allocate the EUE across the time steps. Through this adjustment, we are left with only inefficiencies related to the first two effects, which are relevant to investigation in this paper. The third is out of scope as it relates to how to incentivize effective integration of demand-side flexibility under a price cap. Hence, in this paper, we assume that another scheme has successfully integrated demand-side flexibility to effectively distribute unserved energy.

### v2

@@ II-A4 Market clearing conditions
Finally, agent-specific decisions must comply with system-wide constraints. In particular, market-clearing conditions ensure a balance between supply and demand for both energy and CMs, and a constraint on the annual expected emissions ensures that the CO2 emissions from electricity generation do not exceed the cap. The dual variables of these constraints, $\lambda^{E}_{\omega,t}$, $\lambda^{C}$, and $\lambda^{\text{CT}}$, can be interpreted as market prices (weighted with scenario probabilities and time interval durations).
@@
$=d_{\omega t}^{\text{fix}}-d_{\omega t}^{\text{flex}}\quad\forall t,\omega;\qquad\hskip 0.0pt(\lambda^{E}_{\omega t})$ (8a)
$\sum_{g}c^{\text{CM}}_{g}+\sum_{s}c^{\text{CM}}_{s}=d^{\text{CM}};\qquad\hskip 0.0pt(\lambda^{\text{CM}})$ (8b)
$\sum_{g\omega t}w_{\omega t}\delta_{\omega}q_{\omega t,g}\text{EF}_{g}\leq\bar{E}^{\text{CT}};\qquad\hskip 0.0pt(\lambda^{\text{CT}})$ (8c)
@@ II-B Solution strategy
The market players’ problems are convex quadratic with linear constraints, making their Karush–Kuhn–Tucker (KKT) conditions necessary and sufficient for optimality [mokhtar_s_bazaraa_nonlinear_2005]. By combining KKT conditions and applying the integrability principle of variational inequalities [steven_a_gabriel_complementarity_2013], the solution to the central planner’s convex quadratic optimization problem equates to the equilibrium solution [steven_a_gabriel_complementarity_2013].
@@ II-C Capacity accreditation
We estimate capacity credits using a marginal accreditation approach, supported by academic literature and increasingly adopted in industry [zachary_integration_2022]. Credits are calculated for each generator $g\in\mathcal{G}$ and storage unit $s\in\mathcal{S}$, with storage evaluated across various durations $\zeta\in\mathcal{Z}$. Thus, the resource set for accreditation is $\mathcal{R}=\mathcal{G}\cup(\mathcal{S}\times\mathcal{Z})$.
The accreditation methodology builds upon the structure presented by Wang et al. [wang_crediting_2022] to consider multiple storage durations. The methodology is similar to NESO’s marginal EFC approach as it iteratively perturbs a baseline portfolio [national_energy_system_operator_storage_2024]. [^4] Our accreditation approach is closely aligned with the NESO EFC framework. However, rather than iteratively adjusting the firm capacity to find the quantity of perfectly firm capacity that restores the reliability metric to a target level, we directly evaluate the change in the specified reliability metric (EUE) resulting from a small perturbation of capacity. This approach is aligned with [zachary_integration_2022] and avoids the need for iterative adjustments while preserving the marginal, risk-based interpretation of capacity credits, which is central to NESO’s method. In general, the baseline portfolio can be obtained in two ways. One approach finds a baseline portfolio by solving a welfare-maximizing problem that endogenizes the reliability value by considering the uncapped demand curve and the true WTP of consumers. Another approach finds a baseline portfolio by embedding explicit reliability constraints (e.g., limits on EUE) directly in the planning problem with a price-capped demand curve (as in [da_costa_reliability-constrained_2021]). In this article, we follow the former approach and obtain a baseline portfolio ($\hat{c_{r}}$) that maximizes social welfare in the absence of price cap distortions, integrating the true consumers’ WTP function (where prices can spike up to the VOLL). Dispatching the baseline portfolio under the price-capped market, we determine the baseline $\text{EUE}_{0}$.
Unlike in conventional models with perfectly inelastic demand, the definition of unserved energy is less straightforward when demand is price-elastic. In this work, we follow an approach similar to [kaminski_impact_2021]. Considering all demand with WTP greater than or equal to the price cap as fixed, we record any shortfall of consumption with respect to that demand as unserved [^5] Note that the price cap not only limits the energy price levels during scarcity but also removes part of the flexible demand, which would voluntarily adjust its level based on the energy price. Consistent with our assumption of efficient rationing, we treat this flexible segment as the first portion of demand to be shed once the price cap becomes binding.. The unserved energy $l_{\omega t}$ is given by:
@@
Subsequently, we perturb the baseline portfolio ($\hat{c_{r}}$) by adding a hypothetical perfectly reliable reference generator with small capacity $\epsilon$ and re-dispatching the perturbed portfolio. We use $\text{EUE}_{\text{ref}}$ to note the expected unserved energy obtained from the redispatch of that perturbed portfolio. Similarly, for each resource $r$ (including storage at a fixed duration), we add $\epsilon$ of $r$ to the baseline portfolio, redispatch the perturbed portfolio, and record $\text{EUE}_{r}$.
Following the definition of capacity credit by Wang et al. [wang_crediting_2022], we calculate the marginal capacity credit as the EUE reduction caused by $\epsilon$ additional capacity of $r$, normalized by the EUE reduction caused by $\epsilon$ additional capacity of a perfectly reliable generator. The complete process is presented in Algorithm 1.
@@
  Output: Capacity credit $\text{CC}_{r},\forall r\in R$
@@
  Fix $c_{r}\leftarrow\hat{c}_{r},\forall r\in\mathcal{R}$
   Fix $e_{s}^{\text{init}}\leftarrow\hat{e}_{s}^{\text{init}},\forall s\in\mathcal{S}$
@@
   Compute capacity credit: $\text{CC}_{r}=\frac{\text{EUE}_{0}-\text{EUE}_{r}}{\text{EUE}_{0}-\text{EUE}_{\text{ref}}}$
@@
Note that Algorithm 1 estimates capacity credits that are valid for the baseline portfolio. To capture the portfolio-dependent nature of marginal capacity credits, best practice recommends ELCC and EFC (hyper) surfaces [mantegna_electric_2025, ming_resource_2025]. Depending on the application, capacity credit surfaces are constructed by estimating credits for various portfolios, or the surfaces are explored ad hoc by considering the portfolio obtained with the latest credits. Given the article’s analytical focus, the latter approach is suitable. In our case study, as we show later in Section IV.C, the obtained and baseline portfolios are relatively similar, negating the need for iteratively adjusting the capacity credits.
@@ II-D Capacity market parameters
The capacity market WTP function is constructed as a piece-wise linear curve passing through anchor points specified in the CM rules. The rules determine the coordinates of the anchor points as a function of the system-wide net CONE and capacity target CT. Following [bothwell_crediting_2017], we calculate the capacity target CT as the total accredited capacity of the baseline portfolio. The installed capacities of that portfolio are obtained from the welfare-maximising EOM VOLL run, and denoted $\hat{c}_{g}$ for generators and $(\hat{c}_{s}^{P},\hat{\zeta}_{s})$ for storage (power and duration):
$\text{CT}=\sum_{g}\text{CC}_{g}\,\hat{c}_{g}\;+\;\sum_{s}f^{\text{CC}}(\hat{\zeta}_{s})\,\hat{c}_{s}^{P}.$ (11)
This ensures that the capacity target is consistent with the baseline portfolio used for marginal accreditation, with storage credits evaluated at the equilibrium duration $\hat{\zeta}_{s}$.
Net CONE is defined as the difference between a technology’s annualised fixed and operational costs and its expected net revenues from the energy market, including ancillary services where applicable. For any resource $r$, let $\text{NC}^{\text{uc}}_{r}$ denote the net CONE per installed MW calculated when the baseline portfolio is dispatched under the price cap. For generators and storage, this is given by:
$\text{NC}^{\text{uc}}_{g}=\Bigg[\hat{c}_{g}\left(\widetilde{I}_{g}+I_{g}^{F}\right)-\sum_{\omega,t}w_{\omega t}\,\delta_{\omega}\hat{q}_{\omega tg}\left(\hat{\lambda}^{E}_{\omega t}\right$.
@@
$\left.-C_{g}^{V}-\hat{\lambda}^{\text{CT}}\text{EF}_{g}\right)\Bigg]\frac{1}{\hat{c}_{g}},\quad\forall g\in\mathcal{G},$ (12a)
$\text{NC}^{\text{uc}}_{s}=\Bigg[\hat{c}_{s}^{P}\left(\widetilde{I}_{s}^{P}+I_{s}^{P.F}\right)+\hat{c}_{s}^{E}\left(\widetilde{I}_{s}^{E}+I_{s}^{E.F}\right)$
$-\sum_{\omega,t}w_{\omega t}\,\delta_{\omega}\left(\hat{q}^{\text{dis}}_{\omega ts}-\hat{q}^{\text{ch}}_{\omega ts}\right)\hat{\lambda}^{E}_{\omega t}\Bigg]\frac{1}{\hat{c}_{s}^{P}},\quad\forall s\in\mathcal{S}.$ (12b)
Traditionally, CMs estimate net CONE for a reference technology $\text{NC}_{\text{ref}}$, typically a flexible conventional generator whose accredited capacity is based on its availability. We follow industry practice and use CCGT as the reference generator, assuming, for demonstration purposes, its perfect availability and flexibility.
From an economic perspective, a well-designed CM with a single clearing price should ensure long-run cost recovery. This means that the capacity price per accredited MW should compensate for the missing revenues created by the price cap [zachary_integration_2022, zuo_revisiting_2025]. We assume that a reference generator with a known capacity credit is actively present in the mix. In our setting, this reference generator is CCGT-CCS, which is taken to be perfectly available and flexible so that its accredited capacity equals its installed capacity (i.e., its capacity credit is 1). The long-run break-even condition for this reference generator implies that the capacity market price $\hat{\lambda}^{\text{CM}}$ equals its net CONE per installed MW, $\hat{\lambda}^{\text{CM}}=\text{NC}_{\text{ref}}$. The long-run break even condition for technology $r$ provides us with the missing money-based capacity credit $\text{CC}_{r}^{\text{MM}}$:
$\hat{\lambda}^{\text{CM}}=\frac{\text{NC}^{\text{uc}}_{r}}{\text{CC}^{\text{MM}}_{r}}\;\;\xRightarrow{\hat{\lambda}^{\text{CM}}=\text{NC}_{\text{ref}}}\;\;$
$\text{CC}_{r}^{\text{MM}}=\frac{\text{NC}^{\text{uc}}_{r}}{\text{NC}_{\text{ref}}}\quad\{\forall r\in\mathcal{R}\mid\hat{c}_{r}>0\},$ (13)
The missing money-based capacity credit is the level that would allow each technology $r$ in the baseline portfolio to recover the money that it is missing due to price cap distortions on short-term dispatch and price signals. Therefore, comparing marginal EFCs to $\text{CC}^{\text{MM}}$ indicates whether the capacity credits are at the level required to restore long-run cost recovery of the baseline capacity mix. It is worth noting, though, that the baseline capacity mix does not necessarily yield the maximum social welfare because price caps distort incentives for its optimal dispatch.
Ref. [zachary_integration_2022] proves that marginal EFCs are optimal under multiple assumptions about the power system structure and its operations. In particular, [zachary_integration_2022] assumes that (1) storage and other flexible resources are dispatched to minimise expected unserved energy; (2) storage can be fully recharged between scarcity events, which are assumed to be well separated chronologically; and (3) risk metrics behave smoothly under marginal variations in capacity. In model $M$, where demand is elastic up to the price cap and storage dispatch is optimised for energy price arbitrage, these conditions may not hold. [^6] Solution to model $M$, while maximising the market surplus from dispatch, may not be EUE minimising. In our framework, this assumption may not hold as the model will prefer to shed load rather than charge storage at prices higher than $\text{PC}\cdot\eta_{s}^{\text{ch}}\cdot\eta_{s}^{\text{dis}}$. With price-elastic demand, the system can reduce flexible load up to this threshold, creating spare capacity to charge storage. An EUE-minimising schedule would therefore continue to charge storage or curtail flexible demand beyond the level implied by consumers’ willingness to pay, whereas the market surplus-maximising model will not. This behaviour appears in cases with low emission caps, where the effective marginal cost of gas generation becomes high. Additionally, unlike [zachary_integration_2022], which assumes fixed demand, our model incorporates price-elastic demand, a setting that has not yet been examined in the context of EFC optimality with storage.
@@ II-E Experiment design
For our simulations, we use the model presented in II-A in two distinct market structures: an energy-only market (EOM), where CM-related variables and constraints are omitted, and a combined energy and CM (E+CM). Each structure can be run with different energy demand functions, allowing the price to spike to VOLL or be capped at the price cap, set at an administratively defined level significantly lower than VOLL. When prices are allowed to spike to VOLL, the amount of the price-elastic demand is $D^{\text{flex}}_{\omega,t}$. With the price cap, the width of the linear part of the demand curve reduces by $D^{\text{flex}}_{\omega,t}\cdot\frac{\text{VOLL}-\text{PC}}{\text{VOLL}}$ and the flat part of the demand curve increases in width by the same amount.
We simulate three main cases: EOM VOLL is an energy-only market with prices rising to VOLL; EOM PC is an energy-only market with WTP capped at price cap; E+CM is an energy+capacity market with WTP for energy capped at price cap. When the price cap is applied, the welfare-maximizing mix is no longer financially feasible due to limited price spikes during scarcity, which curb the cost recovery. To identify financially viable capacity mixes in the presence of a cap on energy prices, we solve the E+CM and EOM PC models, which include and lack a capacity market, respectively.
For comparative and accreditation purposes, we also run a case denoted EOM PC opt. mix, where the price is capped, but the capacity mix is fixed to the welfare-maximising mix from EOM VOLL (also called baseline portfolio). The EOM PC opt. mix model provides us with the baseline level for $\text{EUE}_{o}$, which is used to calculate capacity credits.
@@ II-F Performance metrics
We use two primary performance metrics to evaluate the effectiveness of the E+CM model. The first metric is the social welfare (SW), which is calculated as the difference between the consumer benefit (6a) and capital and operating costs. By construction, the SW of the EOM VOLL run will be the highest because in that case, the objective function of the optimisation problem is identical to the SW metric. In the other runs, the SW loss measures the difference in SW between that run and the EOM VOLL run. The second metric is EUE, as defined earlier in (10). We also discuss how different the E+CM and welfare-maximising capacity mixes are.
@@
In this methodological framework, three factors can contribute to social welfare inefficiencies in the E+CM market. First, the shape of the demand curve for the capacity market is administratively set [zhao_constructing_2018]. Even if the shape is correct, calibrating CM parameters is challenging, and estimation errors in capacity credits or net-CONEs can impact the results. Second, imposing price caps narrows price spreads, disrupting storage operational patterns. Third, the unserved demand may be inefficiently allocated between periods as the model assigns a constant value of price cap to every MWh unserved. We evaluate SW according to the true WTP function, following Joskow and Tirole [joskow_reliability_2007] in assuming efficient (non-price) rationing so that customers with the highest WTP are prioritised, but this means the average per-MWh cost of unserved energy rises with the amount unserved in that period. Under the alternative assumption of random rationing, the welfare cost of every MWh unserved would be equal, but the total cost of any given reduction in load would be higher than if it was targeted at the lower-WTP customers.
@@
The latter two effects above relate solely to the introduction of a price cap, and can be assessed by running the EOM PC case with optimal mix obtained from EOM VOLL (EOM PC opt. mix. mentioned earlier). To isolate these effects, we need to obtain the optimal true WTP-aware discharge schedule for the storage and remove the effects of true WTP inobservability on the EUE distribution, i.e., isolating the third effect. First, we fix the generation and charging decisions resulting from the original run. Subsequently, we replace the price-capped WTP with the true WTP, allowing it to reach the VOLL. Finally, we redispatch the mix, with only the discharging variables unfixed, which will now be scheduled to maximise true SW. Because the charging decisions of the storage are fixed, the schedule will only change discharging to better allocate the EUE across the time steps. Through this adjustment, we are left with only inefficiencies related to the first two effects, which are relevant to investigation in this paper. The third is out of scope as it relates to how to incentivize effective integration of demand-side flexibility under a price cap. Hence, in this paper, we assume that another scheme has successfully integrated demand-side flexibility to effectively distribute unserved energy.